OpenStudy (anonymous):

. one number is 3 times as large as the other and the difference between their reciprocals is 1/6. Find the two number.

OpenStudy (anonymous):

1/x -1/(3x) =1/6 or 2/(3x) =1/6 or x=6*2/3 =4 hence the numbers are 4,12